OpenStudy (anonymous):

suppose that y varies directly with x and at the same time varies inversely with Z, and y=12 when x=6 and z=-2. write an equation that models this relationship.

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

you there @perl

OpenStudy (anonymous):

@jim_thompson5910 @mathstudent55 @mathmatition @dan815 @iambatman @Miracrown please help with this question and a few others!!

OpenStudy (perl):

suppose that y varies directly with x and at the same time varies inversely with Z, and y=12 when x=6 and z=-2. write an equation that models this relationship. So the general equation we use to model this is y = k*x / z Your task is to find k , by plugging in values.

OpenStudy (anonymous):

so y=k * 6/-2?

OpenStudy (perl):

you are also given y

OpenStudy (anonymous):

sorry! 12=k*6/-2? 12=k*-3?

OpenStudy (perl):

that looks correct

OpenStudy (anonymous):

well, the 4 answers I am given are y=-4x/z y=4x/z, y=-4z/x and y=4z/x